# Super-Tonks-Girardeau quench of dipolar bosons in a one-dimensional optical lattice

### v1

@@ IV.1 Evaporation mechanism
Figure 3: Dynamics of the reduced two-body density $\rho^{(2)}(x,x^{\prime})$ post quench for $N=5$ bosons and $M=10$ orbitals in $S=5$ sites (unit filling) with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.45E_{r}$ (g)-(i) $g_{d}=-0.82E_{r}$, (j)-(l) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
To clearly explore the many-body features in the sTG quench, we present snapshots of the reduced one-body density $\rho^{(1)}(x,x^{\prime},t)$ and two-body density $\rho^{(2)}(x,x^{\prime},t)$ for some chosen times in Fig. 2 and Fig. 3, respectively. For the TG state at $g_{d}=0.0E_{r}$, the one-body density is perfectly diagonal, indicating only self-correlation within each Mott peak. The two-body density, exhibits a completely depleted region along the diagonal – termed “correlation hole” – which is a manifestation of fermion-like exclusion stemming from the strong repulsion that makes the bosons act as hard-core particles. As already seen from the density, the TG to sTG quench at zero dipolar interactions [Figs. 2(a)-(c) and 3(a)-(c)] is stable for very long times, and both the diagonal shape of $\rho^{(1)}$ and the correlation hole of $\rho^{(2)}$ persist at all probed times. However, when the dipolar interactions are increased such that an eventual instability appears [Figs. 2(d)-(f) and 3(d)-(f)], the diagonal (off-diagonal) pattern in $\rho^{(1)}$ ($\rho^{(2)}$) is maintained only until a critical time, which for the chosen point at $g_{d}=-0.45E_{r}$ is approximately 260 $\mu$s. After that, a rather sudden collapse and eventual evaporation of all correlations occur. Further increasing the interactions, e.g. to $g_{d}=-0.82E_{r}$, leads to a phase transition from the TG Mott state to a progressively more localized cluster state, which is apparent also in the correlation functions [Figs. 2(g)-(i) and 3(g)-(i)]. Five bright spots in $\rho^{(1)}$ distinguish the existence of five strongly interacting bosons at $t=0$. After the quench, at a very small time ($t=5.94\mu$s), the central atom loses its identity, whereas two atoms on each side of it retain their identity although a bit diminished. At $t=29.52$ $\mu$s, we observe complete disruption. However, four very faded traces are visible across the diagonal. Interestingly, the two-body density exhibits correlation across the entire system, even though the density seems localized on the left-hand side. Upon quenching, it spreads and becomes uniform across the whole diagonal and gradually evaporates. For $g_{d}=-2.02E_{r}$, deep in the cluster state, five bright central spots in $\rho^{(1)}$ and a central correlation region in $\rho^{(2)}$ again configure the existence of five bosons. After the quench, they follow the same evaporation process but in a much faster time scale. We remark that the evaporation process is not “macroscopic”; not all the five bosons evaporate simultaneously. It is rather microscopic; the outermost bosons evaporate first and are followed by the rest of the central atoms.
IV.2 Emergence of statistical relaxation and prethermalization
The onset of thermalization in an isolated quantum system with a finite number of interacting particles is an established issue in theoretical physics [69]. The necessary condition for thermalization is statistical relaxation, whereupon various observables settle to some kind of equilibrium. The emergence of prethermalization is characterized by the establishment of a quasi-stationary state at intermediate time scales, followed by relaxation to a stationary state at much longer time scales. Physical origin of prethermalization is still elusive and primarily supposed to be related to integrability modified by external perturbation or quasi-integrability. The time scale of prethermalization in complex many-body systems is crucially determined by the range of interaction and particle number [70]. In the present work, as the sTG phase demonstrates strong many-body correlation and its stability depends on the strength of weak dipolar interactions, we utilize the quenched state to understand the relaxation process, its time scale and the possibility of induce prethermalization.
Many-body information entropy measures such as $S_{\vec{n}}(t)$ defined by Eq. (6) are generally considered as the ideal quantities to study statistical relaxation [68, 67]. To investigate statistical relaxation of the post quench state, we then plot the time evolution of $S_{\vec{n}}$ for different choices of long-range interactions in Fig. 4. As the entropy evolution is very closely related to the occupation in different orbitals and dynamical fragmentation, we also plot the dynamics of the natural occupations in Fig. 5.
For the TG gas without long-range interactions, whose ground state is a five-fold fragmented Mott state, the lowest five orbitals exhibit equal population each of $0.21/N=$ 0.2. Throughout the dynamics, as shown in Fig. 5(a), there is no change in fragmentation because the state is stable against the quenching procedure. Thus, the occupation information entropy remains constant at the initial value of $S_{\vec{n}}\approx 1.61$ as shown in Fig. 4(a).
Now, it is also interesting to assess this constant value utilizing the prediction from the Gaussian orthogonal ensemble (GOE) of random matrices [71]. For time reversal and rotationally invariant systems, statistical relaxation follows the prediction of GOE which estimates $S_{\vec{n}}^{\mathrm{GOE}}=-\sum_{i=1}^{M}\frac{1}{M}\ln\left[\frac{1}{M}\right]=\ln M$, where $M$ is the significantly contributed number of orbitals exhibiting dynamical fragmentation. For the TG to sTG quench without long-range interactions the GOE prediction is then $S_{\vec{n}}^{\mathrm{GOE}}\approx 1.609$ as the system exhibits five-fold fragmentation throughout the dynamics; thus out of $M=10$ orbitals only $M=5$ orbitals are populated. As we can see from Fig. 4(a), the post-quench entropy dynamics perfectly saturates to the value of the GOE prediction.
For $g_{d}=-0.45E_{r}$, Fig. 5(b) shows that initially the system retains five-fold fragmentation as the long-range interaction is weakly attractive. Upon quenching, the system remains highly stable until time $t\approx 236$ $\mu s$, when suddenly the system breaks its stability and becomes fully fragmented. Thus, initially $M=5$ orbitals contribute in fragmentation as well as in the entropy measure, whereas upon complete fragmentation all the $M=10$ orbitals become equally populated. The corresponding entropy measure in Fig. 4(b), starts with the GOE predicted value $\approx 1.61$ ($M=5$) and then increases and saturates at the GOE estimate with $M=10$ orbitals, i.e. $S_{\vec{n}}^{\mathrm{GOE}}\approx 2.30$, which is concomitant with the complete relaxation process to an evaporated gas state.
For $g_{d}=-0.82E_{r}$ and $g_{d}=-2.02E_{r}$ the system becomes fully fragmented very fast as observed in Fig. 5 (c) and (d) respectively. The corresponding entropy production is shown Figs. 4 (c)-(d). In both cases, the entropy measure starts from the initial value close to the GOE prediction, then steeply increases and at the time of complete fragmentation it finally saturates perfectly to GOE estimate. The steep increase in the occupation entropy reflects an exponential increase in the number of excited states participating in the many-body dynamics. For both cases perfect relaxation is achieved and one single time scale is exhibited, although for larger attractive long-range interactions the relaxation happens very fast.
It is also interesting to examine whether the system with long-range attractive interactions always exhibits the monotonic nonequilibrium dynamics or it is possible that before reaching the true relaxed state, the system may exhibit one or several intermediate time scales and more complex relaxation dynamics. To investigate that, we reduce the lattice depth to $V_{0}\approx 8E_{r}$ and slowly control very weak attractive long-range interaction. Results are presented in Fig. 6 for few selected choices of $g_{d}$. It is clearly seen that initially all entropy measures start from $\approx 1.61$ and eventually saturate to $\approx 2.3$ at time $t=59$ $\mu s$. However, they also exhibit intermediate plateaus, where an almost constant entropy is held for certain time. One can define a time $\tau_{\mathrm{pre}}$, as the time for which the prethermal state survives. This is a well separated time scale from the eventual relaxation time scale. We find that $\tau_{\mathrm{pre}}$ is longer for weaker long-range attractions, $g_{d}$ = $-0.2E_{r}$ and $-0.4E_{r}$. With slow tuning and making the long-range interactions stronger, the lifetime of the prethermal state begins to decrease, it becomes quasi-prethermal and then disappears, finally exhibiting single timescale nonequilibrium dynamics for $g_{d}\geq 1.0E_{r}$. This phenomenology indicates that the long-range interaction strength can act as a controlling parameter to reveal relaxation processes in several steps, through well separated timescales. We also find that with other, suitable choices of lattice depths and quenching strengths it is possible to create similar intermediate prethermal or quasi-prethermal states.
Figure 4: Dynamics of occupation entropy for $N=5$ bosons and $M=10$ orbitals in $S=5$ sites (unit filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.45E_{r}$, (c) $g_{d}=-0.82E_{r}$, (d) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
Figure 5: Dynamics of orbital occupation for $N=5$ bosons and $M=10$ orbitals in 5 sites (unit filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.45E_{r}$, (c) $g_{d}=-0.82E_{r}$, (d) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
Figure 6: Dynamics of occupation entropy for $N=5$ bosons and $M=10$ orbitals in $S=5$ sites (unit filling) with $V_{0}\approx 8E_{r}$, $|g_{0}|=20E_{r}$ and (a) $g_{d}=-0.2E_{r}$, (b) $g_{d}=-0.4E_{r}$, (c) $g_{d}=-0.6E_{r}$, (d) $g_{d}=-0.8E_{r}$.
@@ V Results for Double filling
We now present results for more exotic doubly-filled initial states consisting of $N$=6 bosons in $S=3$ sites. Unless otherwise stated, we keep the magnitude of the contact interactions at $|g_{0}|\approx 81E_{r}$ and probe the same range of dipolar interactions $g_{d}\in[-3.6E_{r},0E_{r}]$ as in the single filling case, showing a handful of representative cases.
@@
Figure 7: Density dynamics post quench for $N=6$ bosons and $M=12$ orbitals in $S=3$ sites (double filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.82E_{r}$, (c) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$. The dotted lines indicate the system boundaries.
Fig. 7 illustrate the post-quench density dynamics for increasing values of DDIs $g_{d}$. For all the values of DDIs probed, we find that the system is very unstable to the sudden quench and the density rapidly evaporates. For $g_{d}=0.0$, the initial state consists of pairs of bosons located at the minima of the optical lattice (we again refer to appendix  C for detail on the initial states). Because of the strong contact repulsion, the initial many-body wave function is already fragmented and the particles at each site are spatially separated in dimers –it is a fermionized Mott phase. Upon quenching, the configuration is stable only for short times of around 6 $\mu$s, comparable to the stability of the unit-filling cluster states, after which evaporation and total fragmentation occur, see Fig. 7(a). This is in stark contrast to the unit filling case, where the TG state at $g_{d}=0.0$ exhibits extreme stability. We thus reveal that the stability of the STG state does not survive multiple filling. When the dipolar interactions are turned on, the initial state becomes more clustered. The cluster state is contended by two consecutive lattice sites while keeping the third one completely empty due to frustration emerging between lattice localization, attractive long range interactions and repulsive contact interactions. However, the cluster state clearly exhibits the presence of six bosons: in the density dynamics six bright jets propagate for short time before losing their stability (Fig. 7(b)). By further increasing the dipolar interactions, we achieve an initial cluster state tightly self-bound around the center of the optical lattice. However, six clear humps in the initial density exhibits six strongly correlated bosons in the fermionized limit. Similarly to the fate of the cluster state in the unit filling case, this state is very unstable to quenching. It survives for only a few $\mu$s, before evaporating by rapidly ejecting particles outward as shown in Fig. 7(c).
@@
Figure 8: Dynamics of the reduced one-body density $\rho^{(1)}(x,x^{\prime})$ post quench for $N=6$ bosons and $M=12$ orbitals in $S=3$ sites (double filling) with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.82E_{r}$, (g)-(i) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
@@
Figure 9: Dynamics of the reduced two-body density $\rho^{(2)}(x,x^{\prime})$ post quench for $N=6$ bosons and $M=12$ orbitals in $S=3$ sites (double filling) with (a)-(c) $g_{d}=0.0E_{r}$, (d)-(f) $g_{d}=-0.82E_{r}$, (g)-(i) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
The mechanism for the evaporation process at different dipolar interaction strengths becomes clearer upon examining the one-body and two-body reduced density matrices visualized in Figs. 8 and 9. For the dimerized initial states at both zero and non-zero $g_{d}$, the bosons clearly exhibit non-zero off-diagonal one-body correlation within each dimer already at time zero [Figs. 8(a),(d)]. This act as a channel that allows to dissipate correlation away from the main diagonal, leading to subsequent evaporation at later times [Figs. 8(b)-(c),(e)-(f)]. This cross-correlation between different particles is not present in the TG Mott state at unit filling. A similar mechanism is at play also for the cluster state achieved at stronger dipolar interactions [Figs. 8(g)], but in this case the evaporation occurs at even faster rate [Figs. 8(h)-(i)].
In the two-body reduced density matrix, the structure of the dimerized state is clearly visible both in the off-diagonal peaks and in the diagonal correlation hole [Figs. 9(a),(d)]. Upon quenching, the correlation hole is rapidly reduced while at the same time off-diagonal correlation evaporates. Note, however, that the initial geometric structure of the initial state is visible at later times after evaporation has taken place [Figs. 9(c),(f)]. For the cluster state, a similar initial structure is present, albeit with a much smaller correlation hole due to the self-confinement of the bosonic particles [Figs. 9(g)]. The sudden quench immediately destroys the correlation hole and induces correlation in the outward-ejected particles [Figs. 8(h)]. The process rapidly culminates into evaporation of the cluster state into very fragmented single-particle states.
Finally, we briefly examine the occupation and entropy behavior of the three representative double-filling states examined thus far. A plot of the occupation entropies is shown in Fig. 10 whereas the evolution of the natural occupations is depicted in Fig. 11. All doubly filled states become completely fragmented very fast – in the order of 10 $\mu s$ – and all the orbitals acquire the same population of $1/12$. The corresponding entropy measures in Fig. 10 (a)-(c), show saturation to GOE estimate $S_{\vec{n}}^{\mathrm{GOE}}\approx 2.48$ at longer time for all three choices of long-range interactions. For strong attractive long-range interactions with $g_{d}=-2.02E_{r}$, relaxation happens in a single timescale and possibility of prethermalization is ruled out. However, for weaker long-range interaction some additional features are exhibited in the entropy dynamics at intermediate timescale, which signifies the possibility for the existence of extremely short-lived quasi-prethermalization behavior before increasing almost linearly towards the GOE estimate.
Figure 10: Dynamics of occupation entropy for $N=6$ bosons and $M=12$ orbitals with (a)-(b) $g_{d}=0.0E_{r}$, (c)-(d) $g_{d}=-0.82E_{r}$, (e)-(f) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.
Figure 11: Dynamics of orbital occupation for $N=6$ bosons and $M=12$ orbitals in three sites (double filling) with (a) $g_{d}=0.0E_{r}$, (b) $g_{d}=-0.82E_{r}$, (c) $g_{d}=-2.02E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 81E_{r}$.

### v2

@@ IV.1 Density
Figure 3: Dynamics of the reduced two-body density $\rho^{(2)}(x,x^{\prime})$ post quench for $N=5$ bosons and $M=15$ orbitals in $S=5$ sites (unit filling) with (a)-(c) $g_{d}=-0.16E_{r}$, (d)-(f) $g_{d}=-0.18E_{r}$, (g)-(i) $g_{d}=-0.19E_{r}$, (j)-(l) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
IV.2 Correlations
To clearly explore the many-body features in the sTG quench, we present snapshots of the reduced one-body density $\rho^{(1)}(x,x^{\prime},t)$ and two-body density $\rho^{(2)}(x,x^{\prime},t)$ for some chosen times in Fig. 2 and Fig. 3, respectively. These quantities illustrate coherence and correlation effects at the one-body and two-body level.
For the TG Mott states with long-range attractions down to $g_{d}=-0.16E_{r}$, the one-body density is perfectly diagonal, indicating only self-correlation within each Mott peak. The two-body density, exhibits a completely depleted region along the diagonal – termed “correlation hole” – which is a manifestation of fermion-like exclusion stemming from the strong repulsion that makes the bosons act as hard-core particles. As already seen from the density, the TG to sTG quench at zero or weak dipolar interactions [Figs. 2(a)-(c) and 3(a)-(c)] is stable for very long times, and both the diagonal shape of $\rho^{(1)}$ and the correlation hole of $\rho^{(2)}$ persist at all probed times.
However, further increasing the dipolar interactions, e.g. to $g_{d}=-0.18E_{r}$, leads to a transition from the TG Mott state to a progressively more localized cluster state, which is apparent also in the correlation functions [Figs. 2(d)-(f) and 3(d)-(f)]. In $\rho^{(1)}$, we can distinguish the existence of coherence across the different sites at $t=0$. After the quench, the off-diagonal coherence progressively decreases and almost disappears at time $t=356$ $\mu$s. The diagonal coherence is also visibly squeezed within each site, indicating a change in the on-site correlation structure. A similar fate is observed for $\rho^{(2)}$, in particular with respect to the correlation hole that disappears. For $gd=-0.19E_{r}$, the correlation with the outer sites has completely vanished from the start. During time evolution, we can see how the diagonal correlation becomes progressively more localized under the influence of the attractive interactions [Figs. 2(g)-(i) and 3(g)-(i)].
For yet stronger dipolar interactions, the initial state is a highly localized droplet with nonzero correlations only within the central site. Upon quenching, the correlations are rapidly squeezed to an even narrower range, both in the one-body and two-body reduced density matrix [Figs. 2(j)-(l) and 3(j)-(k)].
In summary, the above analysis of the post quench correlation dynamics shows that when dipolar interactions are strong enough to perturb the initial state away from the pure TG Mott configuration, a collapse under the sTG quench procedure ultimately occurs. On the contrary, the pure and weakly dipolar TG Mott state does not exhibit intersite correlation and can retain its overall correlation pattern despite the contact interaction quench.
IV.3 Autocorrelation function
The density, one-body, and two-body reduced density matrices analyzed so far indicate that the stability of the initial state is retained only under weak dipolar attractions that minimally perturb the TG Mott configuration. However, these quantities provide only few-body information on the system’s time evolution. It remains unclear whether the apparent stability observed in the density for the sTG state persists at the full many-body level. Is there indeed a threshold in $g_{d}$ below which the sTG state remains unperturbed?
To fully characterize the many-body dynamics, a many-body observable is essential. One such observable is the autocorrelation function $\mathcal{F}(t)$, introduced earlier, which measures the overlap or fidelity between the many-body state at time $t$ and its initial configuration. This function provides a direct indication of any deviation from stability during time evolution.
Figure 4 presents the results for the autocorrelation function in the unit-filling setting. To probe the longest possible times with our computational resources, we show results obtained with $M=12$ orbitals. In Appendix D, we also present results for $M=15$ orbitals, which are nearly indistinguishable from those shown here at shorter times, suggesting that our autocorrelation function is converged within the probed timescales.
From Fig. 4(a), we identify three distinct regimes. The first regime, characterized by high fidelity throughout time evolution, shows that the initial state is only weakly perturbed at the many-body level. This region spans values of $g_{d}$ from 0 to around $-0.1E_{r}$. Surprisingly, this is a much narrower range than the density stability would suggest. In the cuts presented in Fig. 4(b), we even observe that perfect autocorrelation is preserved only in complete absence of dipolar interactions. Even minimal dipolar attractions, such as $g_{d}\approx-0.06E_{r}$, eventually reduce the fidelity to around 0.9 at longer times.
The second regime shows significantly reduced stability during dynamics, marked in orange and red shades in Fig. 4(a), corresponding to intermediate values of $g_{d}$ in the interval $[-0.2E_{r},-0.1E_{r}]$. Remarkably, this interval includes both TG Mott states and cluster states. These states exhibit a rapid fidelity decrease during dynamics, with values reaching between $\mathcal{F}\approx 0.3-0.7$ as shown in Fig. 4(b), indicating that the quench strongly affects the state at the many-body level.
Finally, the third regime features a rapid decrease in the autocorrelation function to near-zero values, primarily involving highly localized cluster states and droplets with $g_{d}<-0.19E_{r}$. The collapse of fidelity to zero reflects a complete reorganization of the many-body state following the quench, driven by strong dipolar attractions and quenched contact interactions.
Figure 4: Dynamics of the autocorrelation function $\mathcal{F}(t)$ post quench for $N=5$ bosons and $M=12$ orbitals in $S=5$ sites (unit filling). (a) Dynamics of $\mathcal{F}(t)$ in the entire range of probed dipolar attractions $g_{d}$ from $0.0$ to $-0.4E_{r}$. (b) Dynamics of $\mathcal{F}(t)$ for a few selected values of $g_{d}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
@@ V Results for Double filling
We now present results for more exotic doubly-filled initial states consisting of $N$=6 bosons in $S=3$ sites. Unless otherwise stated, we keep the magnitude of the contact interactions at $|g_{0}|\approx 2E_{r}$ and probe the same range of dipolar interactions $g_{d}\in[-0.4E_{r},0E_{r}]$ as in the single filling case, showing a handful of representative cases.
@@
Figure 5: Density dynamics post quench for $N=6$ bosons and $M=15$ orbitals in $S=3$ sites (double filling) with (a) $g_{d}=-0.16E_{r}$, (b) $g_{d}=-0.17E_{r}$, (c) $g_{d}=-0.19E_{r}$. (d) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$. The dotted lines indicate the system boundaries.
V.1 Density
Fig. 5 illustrates the post-quench density dynamics for increasing values of dipolar interaction strength $g_{d}$. For all the values probed, we find that the system is rather unstable to the sudden quench and the density rapidly exhibits oscillatory patterns and/or collapse under the attractive forces. For $g_{d}=0.0E_{r}$ – and all the way down to $g_{d}=-0.16E_{r}$ – the initial state consists of three pairs of bosons located at the minima of the optical lattice, i.e. a double-filled Mott state. Upon quenching, the three-peak structure is retained for long times. However, from Fig. 5(a) we can clearly see oscillatory patterns in each peak, signalling that the double-filled Mott state is not robust to the sudden quench perturbation. This is in stark contrast to the unit filling case, where the TG Mott state – despite showing instability at the many-body level – still exhibits extreme stability in the density even down to values of $g_{d}=-0.18E_{r}$. We thus reveal that the stability of the sTG state does not survive multiple filling, even at $g_{d}=0.0$.
When the dipolar attractions are increased, the initial state becomes more clustered. This initially appears at around $g_{d}=-0.17E_{r}$, where the central peak absorbs slightly more density than the two side peaks. This is barely visible in Fig. 5(b), but it can be appreciated in the initial density figure in appendix C. When the initial density only slightly deviates from the double-filled Mott state, the dynamics is essentially the same: the three-peak structure is preserved albeit with small on-site oscillations. By further increasing the magnitude of the dipolar interactions to $g_{d}=-0.19E_{r}$, however, we promptly achieve a much more localized cluster state. This cluster state is even more sensitive to the density perturbations caused by the interaction quench, as we can see from the more pronounced oscillations coupled with a narrowing of the peaks in Fig. 5(c).
Finally, when the long-range interactions reach $g_{d}=-0.2E_{r}$, we achieve an initial single-peak state. Similarly to the fate of the single-site droplet in the unit filling case, this state is quite unstable to quenching. As shown in Fig. 5(d), it rapidly undergoes density oscillations that culminate in a much more compact density squeezed by the strong attraction. We also remark that the time scales for this collapse are much faster, in the order of 100 $\mu$s.
@@ V.1 Density
Figure 6: Dynamics of the reduced one-body density $\rho^{(1)}(x,x^{\prime})$ post quench for $N=6$ bosons and $M=15$ orbitals in $S=3$ sites (double filling) with (a)-(c) $g_{d}=-0.16E_{r}$, (d)-(f) $g_{d}=-0.17E_{r}$, (g)-(i) $g_{d}=-0.19E_{r}$, (j)-(l) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.
@@
Figure 7: Dynamics of the reduced two-body density $\rho^{(2)}(x,x^{\prime})$ post quench for $N=6$ bosons and $M=15$ orbitals in $S=3$ sites (double filling) with (a)-(c) $g_{d}=-0.16E_{r}$, (d)-(f) $g_{d}=-0.17E_{r}$, (g)-(i) $g_{d}=-0.19E_{r}$, (j)-(l) $g_{d}=-0.4E_{r}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$. Note that the colorbar is chosen in a range that makes fainter off-diagonal features more visible.
V.2 Correlations
The mechanism for the collapse process at different dipolar interaction strengths becomes again clearer upon examining the one-body and two-body reduced density matrices visualized in Figs. 6 and 7. For the double-filled Mott initial state (both at zero and non-zero $g_{d}$), the bosons display only diagonal one-body correlation throughout the time evolution, with the correlation becoming progressively more confined under the attraction as time goes on [Figs. 6(a)-(c)]. In the two-body reduced density matrix, the initial state clearly reveals two bosons per site, which are separated by a small correlation hole. However, this correlation hole quickly vanishes under the influence of the attractive forces, merging the two correlation blobs along the diagonal into a single squeezed one [Figs. 7(a)-(c)]. The situation is very similar once we deviate slightly from the homogeneous double-filled case at $g_{d}=-0.17E_{r}$ [Figs. 7(d)-(f)]. However, in this case a faint trace of off-diagonal one-body correlation can be identified in the initial state, which seems to persist for longer times [Figs. 6(d)-(f)].
For more prominent cluster states, e.g. at $g_{d}=-0.19E_{r}$, the off-diagonal one-body coherence at time zero is quickly dissipated after the quench [Figs. 6(g)-(i)], while the diagonal correlation is strongly squeezed. In the two-body reduced density matrix, we observe a similar dissipation of off-diagonal correlation between the two side-peaks, while the correlation hole is rapidly populated and the overall correlation is squeezed to narrower regions in the central peak [Figs. 7(g)-(i)].
Finally, below $g_{d}=-0.2E_{r}$ we enter the droplet regime where only one peak is populated. As already observed for the density, this state is highly susceptible to the quantum quench. Both the one-body [Figs. 6(j)-(l)] and the two-body [Figs. 7(j)-(l)] correlation is rapidly confined to smaller areas as the strong attractions overcome the more spread-out initial state.
V.2.1 Autocorrelation function
The stronger instability with respect to the interaction quench exhibited by the double-filled states can be better quantified by looking at the autocorrelation function. This is shown in Fig. 8 again in two ways: in panel (a) by showing the time evolution of $\mathcal{F}(t)$ across all values of $g_{d}$, and in panel (b) for a few selected cuts. In this case, the high-stability region has completely disappeared and only the two cases of low stability (purple region in Fig. 8(a)) and complete many-body disruption (black region in Fig. 8(a)) remain. In fact, all values of $g_{d}$ between 0 and $-0.19E_{r}$ display a rapid decay to a low saturation value of the fidelity around 0.4 with some beating patterns. For even strong dipolar attraction, the autocorrelation function drops instead all the way down to zero, indicating a complete change of the many-body state.
A comparison between single-filling and double-filling systems highlights the critical role of coherence in the stability of the dynamics. In the single-filling case, the loss of off-diagonal coherences significantly contributes to the decay of the initial state. This effect is particularly evident in the sensitivity of the fidelity to the dipolar interaction strength in the region $g_{d}>-0.2E_{r}$ (see Fig. 4). Conversely, in the double-filling case, the dynamics of the fidelity remains largely unchanged within the same $g_{d}>-0.2E_{r}$, which could be attributed the attractive dipolar interactions suppressing off-diagonal coherences. The difference between the two filling scenarios underscores the importance of coherence in determining the stability of the many-body state across different state configurations.
Figure 8: Dynamics of the autocorrelation function $\mathcal{F}(t)$ post quench for $N=6$ bosons and $M=12$ orbitals in $S=3$ sites (double filling). (a) Dynamics of $\mathcal{F}(t)$ in the entire range of probed dipolar attractions $g_{d}$ from $0.0$ to $-0.4E_{r}$. (b) Dynamics of $\mathcal{F}(t)$ for a few selected values of $g_{d}$. The other parameters are $V_{0}\approx 20E_{r}$ and $|g_{0}|\approx 2E_{r}$.